OpenStudy (anonymous):

solve this for x: 22=2(x+6)

OpenStudy (anonymous):

your clever

OpenStudy (igloo3000):

5

OpenStudy (anonymous):

Persistent in pursuit of the answer. Good trait.

OpenStudy (igloo3000):

of course that is how I roll

myininaya (myininaya):

Please don't just give the answers. At least show explanation. @ilovepanda123 Please don't post the same question twice. If you don't understand something ask questions about it. If you aren't getting the attention for that question that you desire, please post a link to your question in chat. We actually have a new code of conduct you might should see. It is pretty awesome. http://openstudy.com/code-of-conduct

OpenStudy (igloo3000):

oh sorry

OpenStudy (anonymous):

:)

OpenStudy (igloo3000):

22=2(x+6) 22=2x+12 10=2x 5